OpenStudy (anonymous):

What is the weight of an object with a mass of 75.0 kg? (Show formula and work.)

OpenStudy (anonymous):

on what planet?

OpenStudy (anonymous):

earth

OpenStudy (ash2326):

weight =Force=mass*gravity =75*9.8 on earth

OpenStudy (anonymous):

which is 735 what?

OpenStudy (ash2326):

newton

OpenStudy (anonymous):

Thank you

OpenStudy (ash2326):

welcome

OpenStudy (anonymous):

The gravitational constant, G, is given a value of 6.67 × 10-11 m3/kg • s2. Using this G value in the law of universal gravitation found on page 76, what force acts between two objects, 3.0 m apart (d) whose masses are 10 kg and 20 kg? (Show formula and work.)

OpenStudy (anonymous):

do you know the gravitational formula?

OpenStudy (ash2326):

F=G*m1*m2/r^2 F=6.67*10^-11*10*20/(3)^2

OpenStudy (anonymous):

Yeah I'm confused on where to put the numbers.

OpenStudy (ash2326):

see m1 =mass of first object m2=mass of second object here m1=10kg and m2=20kg or vice versa it doesn't matter r is the distance betwwn them here it's 3 meters G=universal gravitational constant

OpenStudy (anonymous):

Ohh mmkay.

OpenStudy (ash2326):

got it?

OpenStudy (anonymous):

So the answer would be 1.482?

OpenStudy (ash2326):

1.482*10^-9 Newton

OpenStudy (anonymous):

Okay. Thank you.

OpenStudy (ash2326):

welcome